The foot of the perpendicular drawn from the origin to the plane is (2,5,7). FInd the equation of the plane.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Equation of plane passing through and perpendicular to oa
a=2i+5j+7k
n=OA=2i+5j+7k
Required plane equation is r.n=a.n
r(2i+5j+7k)=(2i+5j+7k)(2i+5j+7k)
so r(2i+5j+7k)=78
In cartesian form 2x+5y+7z=78
Note:the above notations r,n,oa are vectors